OpenStudy (anonymous):

Express log(2x^2-2y^2+1 + i4xy) in the format u +iv where i is the imaginary number

OpenStudy (anonymous):

and \(x, y \) are real?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

\(\log(z)=\ln(|z|) + i \arg(z)\)

OpenStudy (anonymous):

and of course log is a multi valued function, because \(\arg(z)\) is not unique

OpenStudy (anonymous):

\[z=2x^2-2y^2+1 + i4xy\] \(re(z)=2x^2-2y^2+1\) and \(im(z)=4xy\) so \(|z|=\sqrt{(2x^2-2y^2+1)^2+(4xy)^2}\) some algebra is needed here

OpenStudy (anonymous):

i thought this would turn out to be an nice perfect square or something, but it doesn't i wonder if i am doing something wrong it is just what it is

OpenStudy (anonymous):

Can you find a branch then of log(2*z^2+1) that is analytic at i

OpenStudy (anonymous):

now you are making me think

OpenStudy (anonymous):

My next question is to find the derivative of log(2z^2=1) at i. My answer is to use the chain rule and to substitute i with an answer = -4i. Is this the correct way to go?

OpenStudy (anonymous):

yes to your second question the derivative will be \(\frac{1}{2z^2+1}\times 4z\) but i am still thinking about what branch cut you need, and why you cannot do the usual cut on the negative reals

OpenStudy (anonymous):

I am waiting, I thing you can help me with this.

OpenStudy (anonymous):

i just need to find an example. i think this is straight forward. let me check alfors

OpenStudy (anonymous):

damn not there. ok, do you have a similar example to work from?

OpenStudy (anonymous):

No, not at this time. I am going to post another complex analysis question as a new question. We can leave this one, maybe you will solve it late. I first would like to no where do you get the nice mathematical notations in your response seeing that these brackets/squares and exponentiation signs confuse me at times. Thank you for assisting on this problem

OpenStudy (anonymous):

i am going to keep looking and see if i can find how to do this. i think it is a simple matter of finding the zeros of your inside function and cutting there , but i could be wrong

OpenStudy (anonymous):

maybe we can mimic this

OpenStudy (anonymous):

wrong one, i meant this one, problem 4

OpenStudy (anonymous):

Thank you very much. Would you please check the new problem for me.